OpenStudy (anonymous):

18% and 6% of a sum of money differs by $60. How much is the sum of money?

OpenStudy (jamesj):

Call the amount of money X. How much is 18% of X? It is 0.18X. And how much is 6% of X? It is 0.06X Now, what's the relationship between 0.18X 0.06X and $60 ?

OpenStudy (anonymous):

0.12 x $60

OpenStudy (jamesj):

no

OpenStudy (anonymous):

i dont know D:

OpenStudy (jamesj):

In terms of the algebraic expression we wrote above, what is "18% and 6% of a sum of money". The sum of money is X.

OpenStudy (anonymous):

18% + 6% = sum of money

OpenStudy (jamesj):

0.18X and 0.06X Now, what is the difference between the two of them? 0.18X - 0.06X And what amount are we told that is equal to?

OpenStudy (anonymous):

0.12

OpenStudy (jamesj):

No. 0.18X - 0.06X = 0.12X And what now is the amount of money we are told that difference is equal to?

OpenStudy (anonymous):

$60

OpenStudy (jamesj):

Right, so what is the equation you need to solve?

OpenStudy (anonymous):

it need to solve the sum of money?

OpenStudy (jamesj):

"18% and 6% of a sum of money differs by $60. How much is the sum of money?"

OpenStudy (jamesj):

What is the algebraic expression we are using for "the sum of money"? X What do we know about X?

OpenStudy (jamesj):

We know that 0.18X - 0.06X = $60

OpenStudy (jamesj):

Now solve that equation for X.

OpenStudy (anonymous):

0.12X=$60 60/0.12= X

OpenStudy (jamesj):

Yes, but you can't leave the answer that way. Calculate it out.

OpenStudy (anonymous):

500 :D YES YOUR THE BEST TEACHER ONLINE EVER!!

OpenStudy (jamesj):

Now verify the result to make sure you got it right. What is 18% of X?

OpenStudy (anonymous):

I Don't get it

OpenStudy (anonymous):

18% multiply x

OpenStudy (jamesj):

We want to make sure we haven't made a mistake. To do that, we are now working the problem "forwards" to make sure it is consistent with our solution, which was working the problem "backwards". So working it forwards... what is 18% of $500? $90 What is 6% of $500? $30 Hence what is the difference of 18% of $500 and 6% of $500? It is $90 - $30 = $60 Is that consistent with what we were told?

OpenStudy (anonymous):

YES SIR :D

OpenStudy (jamesj):

Yes it is consistent. Hence we can now say with confidence that $500 is the sum of money and the final answer. I strongly recommend you do this verification step at the end of all these problems. You will catch many errors this way.

OpenStudy (anonymous):

Yes thank you for your recommendation.

OpenStudy (anonymous):

Can i ask you something else please

OpenStudy (jamesj):

if it's a new question, post it on the left. But in principle, yes.

OpenStudy (anonymous):

K Thanks for your help, i don't know what else i should do to give you something back

OpenStudy (jamesj):

don't worry. ok, till later

OpenStudy (anonymous):

k thank you again

OpenStudy (anonymous):

i gave you a medal